Solve this system of equations.

(x + 1)2 + (y + 1)2 = 25,x + y = 5


1
Expert's answer
2022-05-29T17:54:02-0400

Solution

From second equation y = 5 – x

Let’s substitute this into first equation:

(x + 1)2 + (6 – x)2 = 25  =>  2x2 – 10x + 37 = 25  =>  x2 – 5x + 6 = 0  =>  x1 = 2, x2 = 3

Substituting these into second equation y1 = 3, y2 = 2

So there are two solutions of the given system of equations:

x1 = 2, y1 = 3 and x2 = 3, y2 = 2